OpenStudy (anonymous):

Rather not post them in the group chat...haha so here to everybody thats awesome enough to be helping me :) HR Diagrams Questions in commmmmmmment

OpenStudy (anonymous):

HR Diagrams Questions Answer the following questions after you have completed the HR Plot. Use complete sentences and submit this file for grading. 1) How does our sun compare to other stars in terms of brightness and temperature? 2) Explain any patterns or trends displayed by the Group 1 stars. 3) Which group of stars represents the Main Sequence? How do you know this? 4) Group 2 and Group 3 stars represent which stages of development? Describe these groups in terms of their relative age, size, brightness, and temperature. 5) What determines if a star will end its life as a neutron star, a white dwarf, or a black hole?

OpenStudy (anonymous):

As an average weighted by the stellar luminosity function, the sun is somewhat above the average temperature and very much above the average luminosity. Only 4% of main sequence stars are significantly brighter and hotter than the sun. About 6% of main sequence stars are about the same brightness and temperature. About 90% of main sequence stars are significantly dimmer and cooler than the sun. However, most of our galaxy's light (about 88%) comes from spectral type B stars, even though only one star in each thousand stars belongs to this type.

OpenStudy (anonymous):

That was for question #1

OpenStudy (anonymous):

DAVE. I swear u better not post a whole bunch of copy pasted stuff...>.< not too much at least pleaaaaaaase

OpenStudy (anonymous):

I wrote this

OpenStudy (saifoo.khan):

WOW!

OpenStudy (anonymous):

Q.2 The periodic table is an arrangement of all the known elements in order of increasing atomic number. The reason why the elements are arranged as they are in the periodic table is to fit them all, with their widely diverse physical and chemical properties, into a logical pattern. If sodium is placed beneath lithium and not next door to fluorine, and potassium is placed beneath sodium to begin another row - and so on - it is found that the vertical lines of elements are chemically similar. These vertical lines are called GROUPS. Horizontal lines of elements are called PERIODS. A set of D-BLOCK ELEMENTS, sometimes called the transition metals, occurs between Groups 2 and 13; these are also chemically similar to each other. Some Groups exhibit striking similarity between their elements, such as Group 1, and in other Groups the elements are less similar to each other, such as Group 4, but each Group has a common set of characteristics.

OpenStudy (anonymous):

5. The mass of the star is what determines whether or not it will die a neutron star, white dwarf or black hole. A star more than 8 solar masses will end its life as a neutron star or a black hole. If it is of 8-10 solar masses, then it will be a neutron star. If its more than 10 solar masses, it will die as a black hole.

OpenStudy (anonymous):

Q.3 his relationship applies to main-sequence stars in the range 0.1–50 solar masses.[50] The amount of fuel available for nuclear fusion is proportional to the mass of the star. Thus, the lifetime of a star on the main sequence can be estimated by comparing it to solar evolutionary models. The Sun has been a main-sequence star for about 4.5 billion years and it will become a red giant in 6.5 billion years,[51] for a total main sequence lifetime of roughly 1010 years. Hence:[52] \begin{smallmatrix} \tau_{\rm MS}\ \approx \ 10^{10} \text{years} \cdot \left[ \frac{M}{M_{\bigodot}} \right] \cdot \left[ \frac{L_{\bigodot}}{L} \right]\ =\ 10^{10} \text{years} \cdot \left[ \frac{M}{M_{\bigodot}} \right]^{-2.5} \end{smallmatrix}

OpenStudy (anonymous):

That has.. NOTHING to do with question 2

OpenStudy (anonymous):

2) group 1 is the main sequence stars. they all appear on an S-shaped curve on the HR diagram, and their tendency is the lower the temperature, the higher the luminosity

OpenStudy (anonymous):

wowsah so... a scattered...hahahaha xD

OpenStudy (anonymous):

sorry, actually the lower the temperature, the lower the luminosity that luminosity scale is confusing!

OpenStudy (anonymous):

IKR!!!!

OpenStudy (anonymous):

yeaah so 3 is easy, it's group 1 and you know this because all the dots are on that curve and they have that tendency

OpenStudy (anonymous):

yea i already have 2, 3 and kinda 5 done. i'm not positive about da rest ^_^

OpenStudy (anonymous):

5) the mass of the star determines how a star will die. if it's a low mass star, it will die as a white dwarf and then become a nebula again, and if it's a high mass star it will !!SUPERNOVA!! and turn in to a neutron star and if it's even BIGGER THAN THAT it'll supernova and turn into a really bright, really dense neutron star.

OpenStudy (anonymous):

hmmmmmmm alrighty i got everything so far. wow thanks so much everybody for your help so far. no idea what it means to me :) ♥

OpenStudy (anonymous):

Alrighty so now all I believe i need is number....4? yea # 4 >.<

OpenStudy (anonymous):

HI!

OpenStudy (anonymous):

hey :)

OpenStudy (anonymous):

whatever weirdo. gtfo please. kidding if u were kidding.

OpenStudy (anonymous):

I was

OpenStudy (anonymous):

then nvm

OpenStudy (anonymous):

Dudeeeeeeee, did you find number 4?:o

OpenStudy (anonymous):

wow. This is a bit of cheating. Which doesn't take you very far.. Learn it and never forget it. Not just, ask people to give you the answer.. ask people to teach you about what your talking about